Solve the given equation over the interval [0.2pi) sin^2(x)-(square root of 3)sin(x)+1= cos^2(x)
Answer by ewatrrr(24785)


Hi
interval [0 - 2pi)
sin^2(x)-√3sin(x)+1= cos^2(x)
(1-cos^2(x) -√3sin(x)+1= cos^2(x)
√3sin(x)= 2(cos^2(x) - 1)
√3 sin(x)= -2sin^2(x)
-√3 /2 = sin(x)
See (c0sx, sinx) summary